the density of concentrated ammonium hydroxide which contain 28 wt% NH3 is 0.899 g/ml. what volume of this reagent should be dilute to 500ml to make 0.25M NH3?


1
Expert's answer
2021-04-28T02:46:00-0400

СM = n / V

n (NH3) = CM x V = 0.25 x 0.5 = 0.125 mol

n = m / M

M (NH3) = 18 g/mol

m (NH3) = n x M = 0.125 x 18 = 2.25 g


m1 (NH3) = (2.25 x 100) / 28 = 8.04 g

density = m / V

V1 (NH3) = m / d = 8.04 / 0.899 = 8.94 ml